Listening to music has long been thought to enhance intelligence, especially during infancy and childhood. To test whether this is true, a researcher records the number of hours that eight high-performing students listened to music per day for 1 week. The data are listed in the table.

Music Listening Per
Day (in hours)
4.2
4.7
4.9
3.7
4.2
5.6
4.0
4.3

(a) Find the confidence limits at a 95% CI for this one-independent sample. (Round your answers to two decimal places.)
to hours per day
Answer #1

Given Data (n = 8):

4.2,4.7,4.9,3.7,4.2,5.6,4.0,4.3

Step 1: Calculate the Sample Mean (Xˉ)

Xˉ=4.2+4.7+4.9+3.7+4.2+5.6+4.0+4.38=35.68=4.45hours

Step 2: Calculate the Sample Standard Deviation (s)

s=(XiXˉ)2n1

Deviations from mean:

(0.25,0.25,0.45,0.75,0.25,1.15,0.45,0.15)

Squared deviations:

0.0625,0.0625,0.2025,0.5625,0.0625,1.3225,0.2025,0.0225

Sum of squared deviations:

2.5s=2.570.60hours

Step 3: Determine the t-Score (95% CI, df = 7)

From t-table:

t0.025,72.365

Step 4: Compute the Margin of Error (ME)

ME=t×(sn)=2.365×(0.608)0.50hours

Step 5: Calculate the 95% Confidence Interval

CI=Xˉ±ME=4.45±0.50=(3.95,4.95)hours

Answer:

The 95% confidence interval for music listening hours per day is 3.95 to 4.95 hours.


answered by: anonymous
Answer #2


The statistic software output for this problem is :

A 95% CI is :

3.95 to 4.95 hours per day